In this problem you will predict the pH of a buffer solution and then predict the new pH after you add NaOH or HCl. Write your answers to three decimal places (X.XXX). The Ka of HC2H3O2 is 1.8×10−51.8×10−5.

1) Calculate the pH of a buffer made from mixing 11.011.0 mL of 0.080.08 M NaC2H3O2 and 9.29.2 mL of 0.080.08 M HC2H3O2.

2) Calculate the pH of the buffer when 5.25.2 mL of 0.0090.009 M NaOH is added to the buffer prepared in part 1.


3) Calculate the pH of the buffer when 4.24.2 mL of 0.0080.008 M HCl is added to the buffer prepared in part 1.
Answer #1

1)

mmoles of NaC2H3O2 = 11 x 0.080 = 0.88

mmoles of HC2H3O2 = 9.2 x 0.080 = 0.736

pH = pKa + log [salt / acid]

     = 4.74 + log [0.88 / 0.736]

pH = 4.823

2)

mmoles of NaOH = 5.2 x 0.0090 = 0.0468

pH = pKa + log [salt + C / acid - C]

     = 4.745 + log [0.88 + 0.0468 / 0.736 - 0.0468]

pH = 4.874

3)

mmoles of HCl = 4.2 x 0.008 = 0.0336

pH = 4.745 + log [0.88 - 0.0336 / 0.736 + 0.0336]

pH = 4.786